Five years ago you took out a​ 5/1 adjustable rate mortgage and the​ five-year fixed rate period has just expired. The loan was originally for $ 310 comma 000 with 360 payments at 4.4 % ​APR, compounded monthly. a. Now that you have made 60 ​payments, what is the remaining balance on the​ loan? b. If the interest rate increases by 0.9 %​, to 5.3 % ​APR, compounded​ monthly, what will be your new​ payments? a. Now that you have made 60 ​payments, what is the remaining balance on the​ loan? The remaining balance on the loan is ​$ nothing. ​(Round to the nearest​ cent.)
Answer #1

1.
=FV(4.4%/12,60,-PMT(4.4%/12,12*30,310000),-310000)=$282,158.96

2.
=PMT(5.3%/12,360-60,-282158.96)=$1,699.16